Find the total amount due for the simple interest loan. (Round your answer to the nearest cent.)

$6100 at 5.3% for 4 years 9 months.
Answer #1

Time = 4 years 9 months = 17/4

Interest = Principal * Rate * time/100 = 6100*5.3%*19/4 = 1535.68

Total Amount = Principal + interest = 6100 + 1535.68 = 7635.68

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well